# ToXCL: A Unified Framework for Toxic Speech Detection and Explanation

### v1

@@ Abstract
[…]
@@ 4.4 ToXCL Performance
Flan-T5 + CLH + TG + TF 81.53normal-$\dagger$ 78.19normal-$\dagger$ 66.49 69.11 67.52 72.14 90.09normal-$\dagger$ 90.08normal-$\dagger$ 47.85 68.93 68.16 85.58
ToXCL 81.53normal-$\dagger$ 78.19normal-$\dagger$ 68.11normal-$\dagger$ 71.21normal-$\dagger$ 70.27normal-$\dagger$ 77.38normal-$\dagger$ 90.09normal-$\dagger$ 90.08normal-$\dagger$ 49.03normal-$\dagger$ 69.93normal-$\dagger$ 68.85normal-$\dagger$ 86.09normal-$\dagger$
@@ 5.1 Case Study: Effectiveness of ToXCL
Index
Input Post
Ground Truth Explanation(s)
GPT-2 (IHC)
Flan-T5 (SBIC)
ToXCL
approx half muslms today result of inbreeding, perhaps this why they are ""hell bent"" on destroying christian based west? inbreed = low iq.
Muslims are result of inbreeding
Muslims are destructive
Muslims are inbred.
you hate like a jew. you are fixatd on hating white christians like a jew.
Jewish people have hate toward other groups
Jews hate white people
Jews don’t like white poeple
funny how you never hear about white men raping women in black" neighborhoods. i know 2 white women who ’
Blacks rape whites.
Blacks are inferior
White men are rapists.
she’s another low iq hollywood liberal . islam is submission and btw if they ran america
Liberals are stupid.
None
Liberals are low IQ people.
rt. : " voters on obamas #immigration plans : drop dead " " #noamnesty "
Immigrants ruin society
None
None
bruh i love white people too how do i join
Minorities are lesser races.
None
None
An Asian woman just gave me a handjob while fingering my ass Which isnt what i meant when i asked her to give me a ’Hand Roll and Pork Intestines. That’s the problem with having the Chinese Massage place right next to the Chinese takeout place
[’chinese women are sex workers.’, "chinese women don’t understand english."]
asians eat pork
asian women only care about sex
you see, i am a doctor. today i had to save a black guy’s life i told his family: i did everything i could, but unfortunately he survived.
[’black people are useless’, ’black people need to be killed’]
black people are worthless
black people aren’t as good as whites
@BlakkHoodie_562 yeah that niggah played good, lowkey i wanted brazil to win
[’black people are the n word’]
None
niggah is a slang term
A bf and gf; are suppose to laugh; at these hoes; together;. A guy should never have these hoes; laughing; at his girl;
[’degrades women by calling them names’, ’women are called hoes’, ’women are hoes’]
None
women are hoes

### v2

@@ Abstract
Nhat M. Hoang1thanks:   Equal contribution, Xuan Long Do2,311footnotemark: 1, Duc Anh Do1, Duc Anh Vu1, Luu Anh Tuan1 thanks:   Corresponding author. 1Nanyang Technological University, Singapore, 2National University of Singapore, 3Institute for Infocomm Research (I2R), A*STAR {nhat005, ducanh003, ducanh001}@e.ntu.edu.sg xuanlong.do@u.nus.edu, anhtuan.luu@ntu.edu.sg
@@ 4.4 ToXCL Performance
Flan-T5 + CLH + TG + TF 81.53$\dagger$ 78.19$\dagger$ 66.49 69.11 67.52 72.14 90.09$\dagger$ 90.08$\dagger$ 47.85 68.93 68.16 85.58
ToXCL 81.53$\dagger$ 78.19$\dagger$ 68.11$\dagger$ 71.21$\dagger$ 70.27$\dagger$ 77.38$\dagger$ 90.09$\dagger$ 90.08$\dagger$ 49.03$\dagger$ 69.93$\dagger$ 68.85$\dagger$ 86.09$\dagger$
@@ 5.1 Case Study: Effectiveness of ToXCL
Index Input Post Ground Truth Explanation(s) GPT-2 (IHC) Flan-T5 (SBIC) ToXCL
(1) approx half muslms today result of inbreeding, perhaps this why they are ""hell bent"" on destroying christian based west? inbreed = low iq. Muslims are result of inbreeding Muslims are destructive Muslims are inbred.
(2) you hate like a jew. you are fixatd on hating white christians like a jew. Jewish people have hate toward other groups Jews hate white people Jews don’t like white poeple
(3) funny how you never hear about white men raping women in black" neighborhoods. i know 2 white women who ’ Blacks rape whites. Blacks are inferior White men are rapists.
(4) she’s another low iq hollywood liberal . islam is submission and btw if they ran america Liberals are stupid. None Liberals are low IQ people.
(5) rt. : " voters on obamas #immigration plans : drop dead " " #noamnesty " Immigrants ruin society None None
(6) bruh i love white people too how do i join Minorities are lesser races. None None
(7) An Asian woman just gave me a handjob while fingering my ass Which isnt what i meant when i asked her to give me a ’Hand Roll and Pork Intestines. That’s the problem with having the Chinese Massage place right next to the Chinese takeout place [’chinese women are sex workers.’, "chinese women don’t understand english."] asians eat pork asian women only care about sex
(8) you see, i am a doctor. today i had to save a black guy’s life i told his family: i did everything i could, but unfortunately he survived. [’black people are useless’, ’black people need to be killed’] black people are worthless black people aren’t as good as whites
(9) @BlakkHoodie_562 yeah that niggah played good, lowkey i wanted brazil to win [’black people are the n word’] None niggah is a slang term
(10) A bf and gf; are suppose to laugh; at these hoes; together;. A guy should never have these hoes; laughing; at his girl; [’degrades women by calling them names’, ’women are called hoes’, ’women are hoes’] None women are hoes